# Complements of caustics of the real $J_{10}$ singularities

### v3

@@ 2.2. Set-valued invariant and virtual Morse functions
A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of class $\Phi_{1}$ or $\Phi_{3}$ is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, and all the corresponding critical values are different and not equal to 0.
@@
If $f$ is a generic polynomial of class $\Phi_{1}$ or $\Phi_{3}$, then the set $V_{f}\subset{\mathbb{C}}^{3}$ defined by the equation $f(x,y)+z^{2}=0$ is a smooth complex surface that is homotopy equivalent to the wedge of ten two-dimensional spheres (see, for example, [9] and [3]). The homology group $H_{2}(V_{f})$ is generated by vanishing cycles (see [3], [12]) that are defined by a system of non-intersecting paths in ${\mathbb{C}}^{1}$ connecting the non-critical value 0 with all critical values of $f$, see Fig. 1. We choose these paths so that those going to complex conjugate non-real critical values are symmetric about the real axis, and those going to real values lie in the upper half-plane where the imaginary parts are smaller than the absolute values of the imaginary parts of all non-real critical values.
@@
An order of these vanishing cycles can also be canonically defined. Specifically, the cycles that vanish at real critical points are listed in ascending order of the corresponding critical values.
@@
A virtual Morse function associated with a generic Morse polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of class $\Phi_{1}$ or $\Phi_{3}$ is a collection of its topological data consisting of
@@
c) the string of positive Morse indices of all real critical points of the function $f(x,y)+z^{2}$, and
@@
Example. Some four virtual Morse functions with eight real critical points are shown in Fig. 2. Two vertical lines in each table indicate the last element of the corresponding virtual Morse function: in all four cases, the number of negative, positive and non-real critical values are 3, 5, and 2, respectively. The real critical points of these functions are only minima and saddles.
@@
A critical point of a virtual Morse function is any column of its data set as in Fig. 2, i.e. a column of the intersection matrix, the intersection index with the set of real points, and a Morse index or the information that the critical point is non-real.
@@
If there are more than one of pairs of non-real critical values of $f$, then there can be more than one virtual Morse function associated with the same real polynomial, because the choice of a system of paths is not homotopically unique: see Fig. 1.
@@
Elementary virtual surgeries of virtual Morse functions include six transformations of their data, modeling the standard local topological surgeries of the corresponding real Morse polynomials, namely
@@
    specifically virtual transformations within the classes of virtual Morse functions associated with the same real Morse polynomials, that are caused by flips of standard systems of paths going from 0 to non-real critical values (see Fig. 1 and [3], and also Figs. 19–21 of volume 2).
The results of all these virtual surgeries are determined by the data of the initial virtual Morse functions. For a detailed description of these standard flips of data, see §V.8 of [12], the explicit formulas for them are given in the comments of the computer program that performs them, see the link on page 3.2. In particular, attempting to perform the surgery $s1$ or $s2$ over real critical values $v_{i}$ and $v_{i+1}$ begins with examining the intersection index $\langle\Delta_{i},\Delta_{i+1}\rangle$ of the corresponding vanishing cycles. If this index is 0, then surgery $s2$ occurs; if the index is 1, then surgery $s1$ occurs; in all other cases the surgery fails. Similarly, a collision of two complex conjugate critical values at a real point not separated from 0 by other critical values follows scenario $s4$ if the intersection index is 0, scenario $s3$ if the index is $1$ or 1-1, and fails in all other cases; in the second case the sign of the intersection index allows us to predict the Morse indices of the newborn real critical points.
@@
If our deformed singularity is neither parabolic nor simple, then we cannot be sure that the real surgery corresponding to a virtual one can be realized any time when these intersection index conditions are satisfied. For simple singularities this realization is guaranteed by the properness of the so-called Lyashko–Looijenga map (see [8]). For parabolic singularities, it follows from the Jaworski’s work [6], [7], see Proposition 7 below.
Let $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ be a generic Morse polynomial of class $\Phi_{1}$ or $\Phi_{3}$.
@@
An (abstract) virtual Morse function of type $f$ is any collection of data as in Definition 4 (i.e. a matrix, two strings and two numbers) obtained from an arbitrary virtual Morse function associated with $f$ by an arbitrary chain of elementary virtual surgeries.
The formal graph of type $f$ is the graph, whose vertices are all virtual Morse functions of type $f$, and two such vertices are connected by an edge if and only if they can be obtained from each other by an elementary virtual surgery.
The virtual component $S(f)$ of the formal graph of type $f$ is its subgraph, whose vertices are only the virtual Morse functions of type $f$ that can be obtained from virtual Morse functions associated with $f$ via arbitrary chains of virtual surgeries $s2,s4,s5,s6$ and $s7$ from Definition 6 (i.e. all surgeries that do not model the collision of critical points).
@@
Denote the four virtual Morse functions in Fig. 2 by the letters A, B, C, and D, as shown in the diagram $\left|{A\ B\atop C\ D}\right|$. The transition $A\to B$ describes the passage $s4$ of two imaginary critical values between the fifth and sixth real critical values. The transitions $A\to C$, $B\to D$ and $C\to D$ are respectively the surgeries $s2$ at which the second and third, the seventh and eighth, and the fourth and fifth real critical values meet and overtake each other. These four elementary surgeries form a commutative diagram, and the chain of passages $A\to B\to D\to C\to A$ forms a cycle in a virtual component of the formal graph.
@@
For all generic polynomials $f$ of the same class $\Phi_{1}$ or $\Phi_{3}$ the formal graphs of type $f$ are the same.
If two generic Morse polynomials $f$ and $\tilde{f}$ belong to the same connected component of the set of Morse polynomials of class $\Phi_{1}$ or $\Phi_{3}$, then their virtual components, $S(f)$, and S(f~)S($\tilde${f}), are the same.
Proof. The first statement follows from the connectedness of the spaces $\Phi_{1}$ and $\Phi_{3}$: any two real Morse polynomials from such a space can be connected by a path that intersects the variety of non-Morse polynomials at finitely many points, each of which corresponds to standard surgery. The second statement follows immediately from the definitions. $\Box$
@@
The formal graph of a (not necessarily Morse) polynomial ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of class $\Phi_{1}$ or $\Phi_{3}$ is the formal graph of an arbitrary generic Morse polynomial of the same class.
The virtual component of a (not necessarily strictly) Morse polynomial of class $\Phi_{1}$ or $\Phi_{3}$ is the virtual component of an arbitrary generic polynomial from the same connected component of the space of all Morse polynomials of this class.
@@
For any generic Morse polynomial $f$ of class $\Phi_{1}$ or $\Phi_{3}$, a virtual Morse function $\varphi$ associated with it, and a virtual Morse function $\tilde{\varphi}$ connected with $\varphi$ by an edge of the formal graph, there exists a generic Morse polynomial $\tilde{f}$ and a path in the space $\Phi_{1}$ or $\Phi_{3}$ connecting $f$ and $\tilde{f}$ and containing only one non-generic point at which it experiences a standard surgery of the same type as the edge $\varphi,\tilde{\varphi}$].
Proof. In the case of class $\Phi_{3}$, the proof repeats the proof of Proposition 2 in [15] for $X_{9}^{+}$ singularities. The only difference is that it considers the canonical complex versal deformation of $J_{10}$ singularities (having exactly the form (2)) instead of the canonical versal deformation of $X_{9}$. In the case $\Phi_{1}$, we reduce our study to this complex deformation using the substitution $x=e^{\pi i/4}\tilde{x}$. $\Box$
@@
Let $f$ be a generic polynomial of class $\Phi_{1}$ or $\Phi_{3}$, then
a)) every virtual Morse function of type $f$ is associated with a generic real polynomial $\tilde{f}$ of the same class $\Phi_{1}$ or $\Phi_{3}$ as $f$;
b)) every virtual Morse function from the virtual component $S(f)$ is associated with some generic real polynomial from the same connected component as $f$ of the space of Morse polynomials of class $\Phi_{1}$ or $\Phi_{3}$;
@@
The set-valued invariant of a real Morse polynomial $f$ of class $\Phi_{1}$ or $\Phi_{3}$ is the set of virtual Morse functions corresponding to all vertices of the virtual component of $S(f)$.
The Card invariant of such a Morse polynomial $f$ is the cardinality of this set of vertices.
Clearly, the set-valued invariant determines the “passport” invariant (which can be read from the bottom line of any virtual Morse function representing it).
@@
If two Morse polynomials of class $\Phi_{1}$ or $\Phi_{3}$ are associated with the same virtual Morse function, then they either belong to the same connected component of the space of generic polynomials of this class, or they belong to components that are mapped to each other by the involution
@@
Each virtual component of the formal graph of type $\Phi_{1}$ or $\Phi_{3}$ is associated with one or two isotopy components of the space of Morse polynomials of the corresponding class $\Phi_{1}$ or $\Phi_{3}$. $\Box$
@@
An isotopy class of Morse polynomials and the associated virtual component of the formal graph of class $\Phi_{1}$ or $\Phi_{3}$ are called achiral (respectively, chiral) if the involution (4) takes this isotopy class to itself (respectively, to a different component).
Finally, our system of isotopy invariants of Morse polynomials of classes $\Phi_{1}$ and $\Phi_{3}$ consists of the set-valued invariant of Definition 9 and the reflection class in the case of chiral classes. By Proposition 9, this system of invariants separates all isotopy classes.
@@ 2.3. D-graph invariant
Let $f$ be a generic polynomial ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of class $\Phi_{1}$ or $\Phi_{3}$ with only real critical points, in particular all ten of its critical values are real and distinct, and 0 is a non-critical value of $f$. The matrix of intersection indices of vanishing cycles $\Delta_{i}\in H_{2}(V_{f})$ (numbered in the ascending order of the corresponding critical values) can be depicted by its Coxeter-Dynkin graph (see e.g. [3]) with ten ordered vertices corresponding to all critical values of $f$. Namely, if the intersection index $\langle\Delta_{i},\Delta_{j}\rangle$ is positive, then the corresponding vertices $v_{i}$ and $v_{j}$ are connected by $\langle\Delta_{i},\Delta_{j}\rangle$ solid segments; if $\langle\Delta_{i},\Delta_{j}\rangle$ is negative then they are connected by $\langle\Delta_{i},\Delta_{j}\rangle$ dashed segments.
@@
The D-graph of a generic real Morse polynomial $f$ with only real critical points is (the isomorphism class of) the oriented graph with vertices labeled with indices 0, 1 and 2, which is obtained from the Coxeter-Dynkin graph of $f$ by
@@
The D-graph of a real Morse polynomial is determined by an arbitrary virtual Morse function associated with this polynomial. Indeed, the intersection matrix and the Morse indices are the elements of the virtual Morse function, and the orientation of the edges follows from the order of the rows and columns of this matrix determined by the order of the corresponding critical values. In this way also D-graphs of arbitrary abstract virtual Morse functions with only real critical points are well defined.
@@
In the restriction to generic Morse polynomials with only real critical points, the D-graphs form an invariant of isotopy classes of Morse functions. This invariant is equivalent to the set-valued invariant from §2.2.
Proof of this theorem repeats the proof of Theorem 2 of [15].

### v5

@@ 2.2. Set-valued invariant and virtual Morse functions
A polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of type $\Phi_{1}$ or $\Phi_{3}$ is generic if it has only Morse critical points in ${\mathbb{C}}^{2}$, and all the corresponding critical values are different and not equal to 0.
@@
If $f$ is a generic polynomial of type $\Phi_{1}$ or $\Phi_{3}$, then the set $V_{f}\subset{\mathbb{C}}^{3}$ defined by the equation $f(x,y)+z^{2}=0$ is a smooth complex surface that is homotopy equivalent to the wedge of ten two-dimensional spheres (see, for example, [12] and [4]). The homology group $H_{2}(V_{f})$ is generated by vanishing cycles (see [4], [17]) that are defined by a system of non-intersecting paths in ${\mathbb{C}}^{1}$ connecting the non-critical value 0 with all critical values of $f$, see Fig. 1. We choose these paths so that those going to complex conjugate non-real critical values are symmetric about the real axis, and those going to real values lie in the upper half-plane where the imaginary parts are smaller than the absolute values of the imaginary parts of all non-real critical values.
@@
An order of these vanishing cycles can also be canonically defined. In particular, the cycles that vanish at real critical points are listed first in ascending order of the corresponding critical values.
@@
A virtual Morse function associated with a generic Morse polynomial $f:({\mathbb{C}}^{2},{\mathbb{R}}^{2})\to({\mathbb{C}},{\mathbb{R}})$ of type $\Phi_{1}$ or $\Phi_{3}$ is a collection of its topological data consisting of
@@
c) the string of positive Morse indices (i.e., positive inertia indices of the quadratic parts) of all real critical points of the function $f(x,y)+z^{2}$, and
@@
Example. Some four virtual Morse functions with eight real critical points are shown in Fig. 2. Two vertical lines in each table indicate the last element of the corresponding virtual Morse function. Namely, in all four cases, the numbers of negative, positive and non-real critical values are 3, 5, and 2, respectively. The real critical points of these functions are only minima and saddlepoints, see the bottom lines of the tables.
@@
A critical point of a virtual Morse function is any column of its data set as in Fig. 2, i.e., a column of the intersection matrix, the intersection index with the set of real points, and a Morse index or the information that the critical point is non-real.
@@
If a real polynomial $f$ has more than one pair of non-real critical values, then there can be more than one virtual Morse function associated with $f$, because the choice of a proper system of paths is not homotopically unique: see Fig. 1.
@@
Elementary virtual surgeries of virtual Morse functions include six transformations of their data, modeling the standard local topological surgeries of the corresponding generic Morse polynomials, namely
@@
    specifically virtual transformations within the classes of virtual Morse functions associated with the same real Morse polynomials, that are caused by flips of standard systems of paths going from 0 to non-real critical values (see Fig. 1 and also Figs. 19–21 of [4], volume 2).
The results of all these virtual surgeries are determined by the data of the initial virtual Morse functions. For a detailed description of these standard flips of data, see §V.8 of [17]. The explicit formulas for them are given in the comments of the computer program that performs them, see the link on page 3.2. In particular, attempting to perform the surgery $s1$ or $s2$ over real critical values $v_{i}$ and $v_{i+1}$ begins with examining the intersection index $\langle\Delta_{i},\Delta_{i+1}\rangle$ of the corresponding vanishing cycles. If this index is 0, then surgery $s2$ occurs; if the index is 1, then surgery $s1$ occurs; in all other cases the surgery fails. Similarly, a collision of two complex conjugate critical values at a real point not separated from 0 by other critical values follows scenario $s4$ if the intersection index is 0, scenario $s3$ if the index is $1$ or 1-1, and fails in all other cases; in the second case the sign of the intersection index allows us to predict the Morse indices of the newborn real critical points.
@@
If our deformed singularities were neither parabolic nor simple, we could not be sure that the real surgery corresponding to a virtual one could be realized at any time when these intersection index conditions are satisfied. For simple singularities this realization is guaranteed by the properness of the Lyashko–Looijenga map (see [11]). For parabolic singularities, this follows from the P. Jaworski’s work [7], [8], see Proposition 7 below.
Let $f:{\mathbb{R}}^{2}\to{\mathbb{R}}$ be a generic Morse polynomial of type $\Phi_{1}$ or $\Phi_{3}$.
@@
An (abstract) virtual Morse function of type $f$ is any collection of data as in Definition 4 (i.e., a matrix, two strings, and two numbers) obtained from an arbitrary virtual Morse function associated with $f$ via an arbitrary finite chain of elementary virtual surgeries.
The formal graph of type $f$ is the graph, whose vertices correspond to all virtual Morse functions of type $f$, and two such vertices are connected by an edge if and only if the corresponding virtual Morse functions can be obtained from each other by an elementary virtual surgery.
The virtual component $S(f)$ of the formal graph of type $f$ is its subgraph, whose vertices are only the virtual Morse functions of type $f$ that can be obtained from virtual Morse functions associated with $f$ via arbitrary finite chains of virtual surgeries $s2,s4,s5,s6,$ and $s7$ from Definition 6 (i.e., all surgeries that do not model the collision of critical points).
@@
Denote the four virtual Morse functions in Fig. 2 by the letters A, B, C, and D in accordance with the diagram $A\ B\atop C\ D$}. The transition $A\leftrightarrow B$ describes the passage $s4$ of two imaginary critical values through the real axis between the fifth and sixth real critical values. The transitions $A\leftrightarrow C$, $B\leftrightarrow D$ and $C\leftrightarrow D$ are the surgeries $s2$ at which respectively the second and third, the seventh and eighth, and the fourth and fifth real critical values meet and overtake each other. These four elementary surgeries form a commutative diagram, and the chain of passages $A\to B\to D\to C\to A$ forms a cycle in a virtual component of the formal graph.
@@
For all generic polynomials $f$ of the same type $\Phi_{1}$ or $\Phi_{3}$, the formal graph of type $f$ is the same.
If two generic Morse polynomials $f$ and $\tilde{f}$ belong to the same connected component of the set of Morse polynomials of type $\Phi_{1}$ or $\Phi_{3}$, then their virtual components $S(f)$ and S(f~)S($\tilde${f}) are the same.
Proof. The first statement follows from the connectedness of the spaces $\Phi_{1}$ and $\Phi_{3}$. Any two real Morse polynomials from such a space can be connected by a path that intersects the variety of non-Morse polynomials at finitely many points, each of which corresponds to a standard surgery. The second statement follows immediately from the definitions. $\Box$
@@
The formal graph of a (not necessarily Morse) polynomial ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of type $\Phi_{1}$ or $\Phi_{3}$ is the formal graph of an arbitrary generic Morse polynomial of the same type.
The virtual component of a (not necessarily strictly) Morse polynomial of type $\Phi_{1}$ or $\Phi_{3}$ is the virtual component of an arbitrary generic polynomial from the same connected component of the space of all Morse polynomials of this type.
@@
For any generic Morse polynomial $f$ of type $\Phi_{1}$ or $\Phi_{3}$, a virtual Morse function $\varphi$ associated with it, and a virtual Morse function $\tilde{\varphi}\neq\varphi$ connected with $\varphi$ by an edge of the formal graph not of type $s7$, there exists a generic Morse polynomial $\tilde{f}$ associated with $\tilde{\varphi}$ and a path in the space $\Phi_{1}$ or $\Phi_{3}$ connecting $f$ and $\tilde{f}$ and containing only one non-generic point at which it experiences a standard surgery of the same type as the edge $\varphi,\tilde{\varphi}$].
Proof. For type $\Phi_{3}$ polynomials, the proof repeats the proof of Proposition 2 in [20] for $X_{9}^{+}$ singularities. The only difference is that it considers the canonical versal deformation of the complex $J_{10}$ singularities, which has exactly the form (2) with complex parameters and the condition $\gamma\neq\pm 1$ instead of the canonical versal deformation of $X_{9}$. For type $\Phi_{1}$, we first embed the deformation (1) into the complex $J_{10}$ deformation as described in §1.3. $\Box$
@@
Let $f$ be a generic polynomial of type $\Phi_{1}$ or $\Phi_{3}$, then
a)) every virtual Morse function of type $f$ is associated with a generic real polynomial $\tilde{f}$ of the same type $\Phi_{1}$ or $\Phi_{3}$ as $f$;
b)) every virtual Morse function from the virtual component $S(f)$ is associated with some generic real polynomial from the same connected component as $f$ of the space of Morse polynomials of type $\Phi_{1}$ or $\Phi_{3}$. $\Box$
@@
The set-valued invariant of a real Morse polynomial $f$ of type $\Phi_{1}$ or $\Phi_{3}$ is the set of virtual Morse functions corresponding to all vertices of the virtual component $S(f)$. The invariant Card of such a Morse polynomial $f$ is the cardinality of this set of vertices.
Clearly, the set-valued invariant determines the “passport” invariant (which can be read from the bottom line of any virtual Morse function).
@@
If two generic Morse polynomials of type $\Phi_{1}$ or $\Phi_{3}$ are associated with the same virtual Morse function, then they either belong to the same connected component of the space of generic polynomials of this type, or they belong to the connected components that are mapped to each other by the involution
@@
Each virtual component of the formal graph of type $\Phi_{1}$ or $\Phi_{3}$ is associated with one or two isotopy classes of Morse polynomials of the corresponding type $\Phi_{1}$ or $\Phi_{3}$. $\Box$
@@
An isotopy class of Morse polynomials and the virtual component associated with this isotopy class are called achiral (respectively, chiral) if the involution (4) takes this isotopy class to itself (respectively, to a different component).
Finally, our system of isotopy invariants of Morse polynomials of types $\Phi_{1}$ and $\Phi_{3}$ consists of the set-valued invariant of Definition 9 and the reflection class in the case of chiral classes. According to Proposition 9, this system of invariants separates all isotopy classes.
Proposition 11.
If a polynomial $f$ of type $\Phi_{1}$ or $\Phi_{3}$ is invariant under the involution ((4)), then the intersection of its $\mathfrak{G}$}-orbit with the space of polynomials ((1)) or ((2)) is also invariant under this involution.
Proof. This intersection point and its image under this involution belong to the same $\mathfrak{G}$}-orbit. According to Proposition 1, this orbit has only one intersection point with the space (1) or (2). $\Box$
@@ 2.3. D-graph invariant
Let $f$ be a generic polynomial ${\mathbb{R}}^{2}\to{\mathbb{R}}$ of type $\Phi_{1}$ or $\Phi_{3}$ with only real critical points, in particular all ten of its critical values are real and distinct, and 0 is a non-critical value of $f$. The matrix of intersection indices of vanishing cycles $\Delta_{i}\in H_{2}(V_{f})$ (numbered in the ascending order of the corresponding critical values) can be depicted by its Coxeter-Dynkin graph (see e.g. [4]) with ten ordered vertices corresponding to all critical values of $f$. Namely, if the intersection index $\langle\Delta_{i},\Delta_{j}\rangle$ is positive, then the corresponding vertices $v_{i}$ and $v_{j}$ are connected by $\langle\Delta_{i},\Delta_{j}\rangle$ solid segments; if $\langle\Delta_{i},\Delta_{j}\rangle$ is negative then they are connected by $\langle\Delta_{i},\Delta_{j}\rangle$ dashed segments.
@@
The D-graph of a generic real Morse polynomial $f$ with only real critical points is (the isomorphism class of) the oriented graph with vertices labeled with indices 0, 1 and 2, which is obtained from the Coxeter-Dynkin graph of $f$ via
@@
The D-graph of a real Morse polynomial is determined by an arbitrary virtual Morse function associated with this polynomial. Indeed, the intersection matrix and the Morse indices are the elements of the virtual Morse function, and the orientation of the edges follows from the order of the rows and columns of the intersection matrix, which is determined by the order of the corresponding critical values. In this way, D-graphs of arbitrary virtual Morse functions with only real critical points are also well-defined.
@@
In the restriction to the space of generic Morse polynomials of class $\Phi_{1}$ or $\Phi_{3}$ with only real critical points, the D-graphs form an invariant of isotopy classes of Morse functions. This invariant is equivalent to the set-valued invariant from §2.2.
Proof of this theorem repeats the proof of Theorem 2 of [20]. $\Box$